### v1

# Collective steering: Tracer-informed dynamics

Abstract
We consider control and inference problems where control protocols and internal dynamics are informed by two types of constraints. Our data consist of i) statistics on the ensemble and ii) trajectories or final disposition of selected tracer particles embedded in the flow. Our aim is i’) to specify a control protocol to realize a flow that meets such constraints or ii’) to recover the internal dynamics that are consistent with such a data set. We analyze these problems in the setting of linear flows and Gaussian distributions. The control cost is taken to be a suitable action integral constrained by either the trajectories of tracer particles or their terminal placements.
1 Introduction
The theme of this work is the control problem of steering a collection of particles/agents that obey a universal time-varying feedback law; the law represents a common control protocol broadcast to all agents in the collective. In this paper, we are interested in both control and inference problems. The data of these problems concern, on the one hand, the collective in the form of marginal distribution constraints and, on the other, the locations or trajectories of specific tracer particles embedded in the flow. Whether we consider inference or control, these tracer particles reveal internal dynamics or constrain a sought-after control protocol.
The practical motivation for this class of problems stems from a range of applications in engineering and physics. To mention a few, this is the case in particle systems steered by externally regulated control potentials. Similarly, in swarm robotics, exogenous control inputs are used to steer the swarm into particular configurations. On the flip side, landmarks in the field of view of a collective may serve as tracer particles’ constraints to guide the motion of the ensemble overflying a known terrain. Indeed, also in fluids, tracer particles may reveal the internal dynamics of the flow.
The present work aims to develop this circle of ideas in the most basic setting of linear flows and Gaussian distributions. Accordingly, we consider linear time-varying feedback control laws. Control costs penalize both the kinetic energy of the ensemble as well as the control gain. The former is physically motivated and underlies the theory of optimal mass transport [1, 2, 3, 4]. The latter quantifies the complexity of control implementation as was eloquently explained by R.W. Brockett [5].
In linear state feedback, when the feedback gain is regulated and represents the control authority, the dynamics are bilinear. Due to this fact, controllability issues are subtle and often overlooked. For instance, while it was widely accepted that in this basic setting, the Liouville equation is controllable under standard conditions [6], a proof of this fact was not available until recently [7] (see also, [8]). At the heart of the technical issue is that solutions to optimal control problems for linear systems may not always be expressible in linear feedback form due to topological obstructions [8, Section IV]. Thus, a main contribution of the present work, besides raising attention to this type of problems where tracer particles may inform control and inference, is to establish the existence of solutions to such problems under suitable regularizations of the cost.
Below, we introduce notation and discuss the general setting in Section 2. In Section 3, we introduce and solve the first of our problems: the problem of collective steering with the information on the running distribution of the collective in conjunction with the endpoint positions of tracers. In Section 4, we formulate and solve a dual problem entailing information on the endpoint distributions of the collective and trajectories of the tracers. Lastly, we present an academic example for each problem in Section 5.
2 Problem setting
Throughout, we consider a collection of particles in transport over the time window $t\in[0,1]$ that are subject to the linear dynamics
$\dot{X}_{t}$ $=u_{t}(X_{t})=K_{t}X_{t}.$ (1)
We assume that $X_{0}\sim\mathcal{N}_{n}(0,\Sigma_{0})$ with $\Sigma_{0}\in\mathrm{Sym}^{+}(n)$, where $\mathcal{N}_{n}(0,\Sigma_{0})$ denotes the $n$-dimensional Gaussian distribution with zero mean and covariance $\Sigma_{0}$, and
${\rm Sym}^{+}(n):=\{\Sigma\in\mathbb{R}^{n\times n}:\Sigma\succ 0,\Sigma=\Sigma^{\top}\},$
i.e., the space of $n\times n$ positive-definite and symmetric matrices. If ${\rm GL}(n)$ denotes the real general linear group of degree $n$, i.e., the multiplicative group of $n\times n$ invertible matrices, then the flow of (1) takes the form
$X_{t}=\Phi_{t}X_{0},$ (2)
where the state-transition matrix $\Phi_{t}\in{\rm GL}(n)$ satisfies
$\dot{\Phi}_{t}=K_{t}\Phi_{t},\ \ \Phi_{0}=I.$ (3)
As is well-known, the space of Gaussian distributions is invariant under the action of linear flows, i.e., flows of the form (2), via the pushforward of measures. Then, with direct computation, one can verify that $X_{t}\sim\mathcal{N}_{n}(0,\Phi_{t}\Sigma_{0}\Phi_{t}^{\top})$.
As outlined in the introduction, the data of the problems we consider in this paper belong to one of two types: i) marginal distributional constraints and ii) trajectories/displacement constraints of tracer particles. Expressed in terms of the state transition matrix $\Phi_{t}$ defining the flow of (1), the first of these takes the form
$\Sigma_{t}=\Phi_{t}\Sigma_{0}\Phi_{t}^{\top},\quad\mbox{ with }t\in\mathcal{T}_{1}\subseteq[0,1],$ (4)
where $\Sigma_{t}\in\mathrm{Sym}^{+}(n)$. We will only be interested in the cases where $\mathcal{T}_{1}=\{0,1\}$ or $\mathcal{T}_{1}=[0,1]$. In the latter case, we shall always assume that the curve $\Sigma_{t\in[0,1]}$ is continuously differentiable.
The second type of constraints, ii), can be succinctly explained as follows. Let $y_{t}^{(i)}$ be the position of the $i$-th tracer particle at the time instant $t\in\mathcal{T}_{2}\subseteq[0,1]$ for $i\in\{1,\ldots m\}$ with $m<n$, and let
$Y_{t}=\begin{bmatrix}y_{t}^{(1)}&\ldots&y_{t}^{(m)}\end{bmatrix}\in\mathbb{R}^{n\times m}$
be the matrix comprising instantaneous information on the tracer positions. Then, the constraint ii) is of the form
$Y_{t}=\Phi_{t}Y_{0},\quad\mbox{ with }t\in\mathcal{T}_{2}\subseteq[0,1].$ (5)
Once again, we will only be interested in the cases where $\mathcal{T}_{2}=\{0,1\}$ or $\mathcal{T}_{2}=[0,1]$, and in the latter case, we shall always assume that the curve $Y_{t\in[0,1]}$ is continuously differentiable and of full column rank.
To wrap up this preliminary section, we now introduce the cost functionals we deal with. The first control cost we consider is the integral of the (expected) kinetic energy
$\frac{1}{2}\int_{0}^{1}\mathbb{E}[\|\dot{X}_{t}\|^{2}]\,\mathrm{d}t,$ (6)
where $\mathbb{E}[\cdot]$ denotes the expectation operator (recall that $X_{0}\sim\mathcal{N}(0,\Sigma_{0})$, i.e., $X_{0}$ is a Gaussian random variable). The functional (6) is the typical cost in optimal mass transport and quantifies the control energy, see, e.g., [2, 4]. Substituting (1) and (2) into (6) while utilizing the properties of the expectation operator and the fact that $\mathbb{E}[X_{0}X_{0}^{\top}]=\Sigma_{0}$, we obtain that the functional (6) can be rewritten as
$\int_{0}^{1}{\rm tr}(K_{t}\Phi_{t}\Sigma_{0}\Phi_{t}^{\top}K_{t}^{\top})\,\mathrm{d}t.$ (7)
Notice that, in terms of $K_{t}$ and $\Phi_{t}$, the cost functional (7) is not jointly convex, and so establishing the existence of solutions of optimal control problems with the cost (7) is non-trivial. Indeed, for some terminal specifications the minimizing (7) leads to a singular state transition $\Phi_{t}$ and, consequently, a non-integrable $K_{t}$ [8, Example 1]. The second cost functional we consider is
$\int_{0}^{1}\mathbb{E}[\|\nabla u(X_{t})\|^{2}]\,\mathrm{d}t,$ (8)
which is a modification of the attention functional introduced by R.W. Brockett in [5] and represents a weighted complexity of implementing the control law $u_{t}$ in (1) (the weighting in the form of the expectation is introduced to ensure a finite integral). From (1) and (2), we see that (8) is equivalent to
$\int_{0}^{1}{\rm tr}(K_{t}K_{t}^{\top})\,\mathrm{d}t.$ (9)
Note that unlike the functional (7), the cost functional (9) is convex in $K_{t}$ and is independent of $\Phi_{t}$. It turns out that utilizing the functional (9) as a regularization term for the cost functional (7) is instrumental in establishing the existence of solutions to one of the problems we consider.
3 Internal Dynamics from Tracer Displacements
In this section, we tackle the first of the problems we propose in this paper. Throughout this section, we assume knowledge about the ensemble of particles in the form of the marginal distributional constraints (4) with $\mathcal{T}_{1}=[0,1]$ and the endpoint conditions on the tracers (5) with $\mathcal{T}_{2}=\{0,1\}$. Our task is to identify the dynamics in (3), effected by a suitable control vector field $u_{t}$ as in (1), that realizes the given data about the flow. We take the functional (7) as our control cost. The problem we investigate can then be formulated as follows.
Problem 1
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=[0,1]$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=\{0,1\}$, find the minimizer of (7), subject to the constraints (3), (4), and (5).
Let us define what shall be viewed as an admissible solution to the previous problem. We say a curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is admissible for Problem 1 if $K_{t\in[0,1]}$ is Lebesgue-integrable and $\Phi_{t\in[0,1]}$ is absolutely continuous, and are such that the differential constraint (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists) as well as constraints (4) and (5), and the value of the functional (7) is finite. The following proposition provides a more explicit characterization of the class of admissible curves.
Proposition 1
A curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is an admissible curve for Problem 1 if and only if the following four conditions are satisfied. First,
$\Phi_{1}\Sigma_{0}\Phi_{1}^{\top}=\Sigma_{1},$ (10)
second, the constraint (5) is satisfied, third, there exists an almost everywhere unique Lebesgue-integrable curve $\Omega_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ such that
$K_{t}=\frac{1}{2}\dot{\Sigma}_{t}\Sigma_{t}^{-1}+\Omega_{t}\Sigma_{t}^{-1},\ \ \mbox{ with }\ \ \Omega_{t}^{\top}=-\Omega_{t},$ (11)
for almost all $t\in[0,1]$, and fourth, the differential constraint (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists).
Proof 3.1.
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an admissible curve for Problem 1. Since $\mathcal{T}_{1}=[0,1]$ and $\Sigma_{t\in\mathcal{T}_{1}}$ is continuously differentiable by assumption, direct differentiation of (4) followed by substitution of (3) shows that $K_{t\in[0,1]}$ must satisfy the Lyapunov equation
$\dot{\Sigma}_{t}=K_{t}\Sigma_{t}+\Sigma_{t}K_{t}^{\top},$ (12)
for almost all $t\in[0,1]$. On the other hand, since $\Sigma_{t}\in\mathrm{Sym}^{+}(n)$, it follows from basic linear algebra that the constraint (12) is satisfied for all $t\in\mathcal{T}_{1}=[0,1]$ if and only if $K_{t}$ is given by (11), where the Lebesgue-integrable curve $\Omega_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ is almost everywhere uniquely defined. Finally, since $1\in\mathcal{T}_{1}=[0,1]$, the curve $\Phi_{t\in[0,1]}$ must be such that $\Phi_{1}$ satisfies (10). The converse implication follows by direct integration of the differential constraint (3) with $K_{t\in[0,1]}$ defined by (11).
Next, we observe that if $(K_{t},\Phi_{t})_{t\in[0,1]}$ is any admissible curve for Problem 1, the functional (7) is given by
$\int_{0}^{1}{\rm tr}(K_{t}\Sigma_{t}K_{t}^{\top})\,\mathrm{d}t,$
where we have substituted the constraint (4) into the functional (7). The advantage of the latter representation of the functional is that it is convex in the free parameter $\Omega_{t}$ (after another substitution from (11)) and is independent of $\Phi_{t}$. The preceding discussion suggests the reformulation of Problem 1 as follows.
Problem 1
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=[0,1]$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=\{0,1\}$, find the minimizer of
$\int_{0}^{1}{\rm tr}(K_{t}\Sigma_{t}K_{t}^{\top})\,\mathrm{d}t$
subject to the constraints (3), (5), (10), and (11).
The next proposition, the proof of which is a direct consequence of Proposition 1 and the discussion preceding the statement of Problem 1, formulates the equivalence between Problem 1 and Problem 1.
Proposition 1.
An admissible curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ solves Problem 1 if and only if it solves Problem 1.
Before we proceed with deriving the necessary conditions of optimality, we state in passing a theorem on the existence of solutions to Problem 1 and, by extension, Problem 1.
Theorem 2.
Suppose at least one admissible curve exists for Problem 1. Then, there exists an absolute minimizer $(K_{t},\Phi_{t})_{t\in[0,1]}$ for Problem 1.
The theorem implies that Problem 1 has an absolute minimizer, which, however, may not be unique. The proof, due to its technical nature, is omitted. Briefly, it consists of verifying that the conditions of [9, Theorem 11.4.vi] hold. Thus, in light of Theorem 2, we can now meaningfully seek a minimizer by solving the necessary conditions of optimality obtained from the classical theory of optimal control [10] as formulated next.
Proposition 3.
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an absolute minimizer of Problem 1, or equivalently Problem 1. Then,
$\dot{\Phi}_{t}$ $=K_{t}\Phi_{t},$ (13a)
with $\Phi_{0}=I$ and $K_{t}$ is given by (11) and
$\Omega_{t}\Sigma_{t}+\Sigma_{t}\Omega_{t}$ $=$
$\frac{1}{2}(\Phi_{t}P_{t}^{\top}+\dot{\Sigma}_{t})\Sigma_{t}-\frac{1}{2}\Sigma_{t}(P_{t}\Phi_{t}^{\top}+\dot{\Sigma}_{t}),$ (13b)
where $\text{where }\hskip 20.0pt\dot{P_{t}}$ $=-{K_{t}}^{\top}P_{t}.$ (13c)
Moreover, if $\mathcal{S}\subset\mathrm{GL}(n)$ is the surface defined by
$\mathcal{S}:=\{\Phi\in\mathrm{GL}(n)~{}|~{}\Phi\Sigma_{0}\Phi^{\top}=\Sigma_{1},\,\Phi Y_{0}=Y_{1}\},$ (13d)
then (13a) and (13c) are subject to the mixed boundary conditions $\Phi_{1}\in\mathcal{S}$ and ${P_{1}}\perp T_{\Phi_{1}}\mathcal{S}$, respectively.
Proof 3.2.
From Proposition 1, the control Hamiltonian corresponding to Problem 1 is
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}K_{t}\Sigma_{t}K_{t}^{\top}+P_{t}^{\top}K_{t}\Phi_{t}\big{)},$
for $K_{t}$ satisfying (11) and where $P_{t}$ is an $n\times n$ Lagrange multiplier (costate) matrix enforcing (3). The first order necessary conditions in $P_{t}$ and $\Phi_{t}$ are such that $\dot{\Phi}_{t}=\partial\mathcal{H}/\partial P_{t}$ and $\dot{P}_{t}=-\partial\mathcal{H}/\partial\Phi_{t}$, respectively. These two conditions yield Equations (13a) and (13c), in that order.
Using the form (11), we rewrite the previous Hamiltonian as
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}\Sigma_{t}^{-1}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}^{\top}\big{)}$
$+{\rm tr}\big{(}\big{(}P_{t}^{\top}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}\Sigma_{t}^{-1}\Phi_{t}\big{)},$
bearing in mind that $\Omega_{t}$ is skew-symmetric. In this case, Pontryagin’s maximum principle is such that $\partial\mathcal{H}/\partial\Omega_{t}=0$. Using that and the fact that
$\frac{\partial}{\partial\Omega_{t}}{\rm tr}\big{(}\Omega_{t}S_{t}\big{)}=\frac{1}{2}\big{(}S_{t}-S_{t}^{\top}\big{)},$
for $S_{t}\in\mathbb{R}^{n\times n}$, the Lyapunov equation (13b) can be obtained.
The terminal condition (10) along with tracers’ condition (5) define the surface $\mathcal{S}$ in (13d). Thus, it is necessary that $\Phi_{1}\in\mathcal{S}$. According to [10, Section 5.1], any admissible variation for the optimal $\Phi_{1}$, $\delta\Phi_{1}$, has to be tangent to $\mathcal{S}$, i.e., $\delta\Phi_{1}\in T_{\Phi_{1}}\mathcal{S}$. We note,
${\rm tr}({P_{1}}^{\top}\delta\Phi_{1})=0,$
for any admissible variation $\delta\Phi_{1}$, cf. [10, Equation 5.1-18], which necessitates that ${P_{1}}\perp T_{\Phi_{1}}\mathcal{S}$.
Any solution to Problem 1 must satisfy the necessary conditions of optimality stated in Proposition 3, constituting a nonlinear two-point boundary value problem. A solution to this problem, which is guaranteed to exist as a consequence of Theorem 2, provides a candidate flow that reconciles the data of Problem 1 with the particle dynamics (1) using a suitable linear feedback law. Numerical results, in addition to visualization of the flow, are presented in Section 5 to illustrate the nature of the solutions to this problem.
4 Internal Dynamics from Tracer Trajectories
This section deals with a counterpart of Problem 1. In it, the data of the problem we consider here is in the form of the distributional constraints (4) with $\mathcal{T}_{1}=\{0,1\}$ and the full trajectories of the tracers (5) with $\mathcal{T}_{2}=[0,1]$. Our task is once again to identify the dynamics in (3). If (7) is taken as the cost of transport, then a candidate problem formulation can be stated as follows.
Problem 4.1.
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find the minimizer of (7), subject to the constraints (3), (4), and (5).
However, this candidate formulation may not admit a solution, in general; there are cases where $\Phi_{t}$ becomes singular at some point in time, see [8, Example 1]. In order to avert this issue, we modify the cost in Problem 4.1 by augmenting (7) with an added “weighted attention” term (9) that penalizes the $L_{2}$-norm of the gain $K_{t}$. Such an augmentation traces back to R.W. Brockett [5], according to whom, “… the easiest control law to implement is a constant input. Anything else requires some attention. The more frequently the control changes, the more effort it takes to implement it. […], it can be argued that the cost of implementation is linked to the rate at which the control changes with changing values of $x$ [state] and $t$ [time].”
For convenience, we restrict ourselves to a spatial attention term as in (8) or (9). Thereby, we state the adaptation of Problem 4.1 as follows.
Problem 4.2.
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find the minimizer of
[…]
for $\varepsilon>0$ and subject to the constraints (3), (4), and (5).
A curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is said to be admissible for Problem 4.2 if $K_{t\in[0,1]}$ is Lebesgue-integrable and $\Phi_{t\in[0,1]}$ is absolutely continuous, and are such that (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists) as well as constraints (4) and (5), and the value of the functional (14) is finite. The following proposition provides a more explicit characterization of the class of admissible curves.
Proposition 4.
Fix $N_{0}\in\mathbb{R}^{(n-m)\times n}$ such that
$N_{0}Y_{0}=0,~{}\mbox{ and }\ \ \mathrm{rank}([Y_{0},\,N_{0}^{\top}])=n.$ (15)
Then, a curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is an admissible curve for Problem 4.2 if and only if the following four conditions hold. First, the constraint (4) is satisfied, second,
$Y_{1}=\Phi_{1}Y_{0},$ (16)
third, there exists an almost everywhere unique Lebesgue integrable curve $R_{t\in[0,1]}\in\mathbb{R}^{n\times(n-m)}$ such that
$K_{t}$ $=M_{t}+R_{t}N_{t},$ (17a)
[…]
for almost all $t\in[0,1]$, and fourth, the constraint (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists).
Proof 4.3.
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an admissible curve for Problem 4.2. Since $\mathcal{T}_{2}=[0,1]$ and $Y_{t\in\mathcal{T}_{2}}$ is continuously differentiable by assumption, direct differentiation of (5) followed by substitution of (3) shows that $K_{t\in[0,1]}$ must satisfy
$\dot{Y}_{t}=K_{t}Y_{t},$ (18)
for almost all $t\in[0,1]$. It follows that
$K_{t}=M_{t}+Z_{t},$ (19)
where $Z_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ is any Lebesgue-integrable curve satisfying $Z_{t}Y_{t}=0$, i.e., the column space of $Y_{t}$ is a subspace of the null space of $Z_{t}$.
To proceed, let $N_{0}\in\mathbb{R}^{(n-m)\times n}$ be any full row-rank matrix such that $N_{0}Y_{0}=0$ and define the curve $N_{t\in[0,1]}$ as the solution to
$\dot{N}_{t}$ $:=-N_{t}M_{t},$ (20)
from the initial condition $N_{0}$, where $M_{t}$ is as in (17b). We claim that $N_{t}$ is of full row rank, and the null space of $N_{t}$ equals the column space of $Y_{t}$ for almost all $t\in[0,1]$. To see this, note that (20) is a linear time-varying system of equations, and so the solution to (20) from the initial condition $N_{0}$ is uniquely given by
$N_{t}$ $=N_{0}\Psi_{t},$
where $\Psi_{t}$ is the fundamental solution matrix associated with the linear system (20). Since $\Psi_{t}$ is invertible for every $t$, it follows that $N_{t}$ is full row rank if and only if $N_{0}$ is full row rank. On the other hand, direct differentiation of the matrix $N_{t}Y_{t}$ followed by a substitution of (17), (19) and (20) yields
$\dot{N}_{t}Y_{t}+N_{t}\dot{Y}_{t}$ $=0,$
implying
$N_{t}Y_{t}=N_{0}Y_{0}=0,\ \ \forall t\in[0,1].$
From the above, it follows that any Lebesgue-integrable curve $Z_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ satisfying the constraint $Z_{t}Y_{t}=0$ must be of the form $Z_{t}=R_{t}N_{t}$, for an almost everywhere unique Lebesgue-integrable curve $R_{t\in[0,1]}\in\mathbb{R}^{n\times(n-m)}$. Finally, since $1\in\mathcal{T}_{2}=[0,1]$, the curve $\Phi_{t\in[0,1]}$ must be such that $\Phi_{1}$ satisfies (16). The converse implication of the proposition follows by direct integration.
Proposition 4 suggests a reformulation of Problem 4.2, replacing condition (5) on the state $\Phi_{t}$ with an explicit constraint on the gain $K_{t}$. The advantage of the new formulation lies in the fact that the search for the optimal curve reduces to searching over a lower dimensional control space, namely, the curve $R_{t\in[0,1]}$.
Problem 4.2
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find the minimizer of (14) for $\varepsilon>0$ and subject to the constraints (3), (4), (16), and (17).
Before we proceed with deriving the necessary conditions of optimality, we state in passing a theorem on the existence of solutions to Problem 4.2.
Theorem 5.
Suppose that there exists at least one admissible curve for Problem 4.2. Then, there exists an absolute minimizer $(K_{t},\Phi_{t})_{t\in[0,1]}$ for Problem 4.2.
Similar to Theorem 2, the proof of the above statement consists of verifying that the conditions of [9, Theorem 11.4.v.i] hold. We now proceed to state the necessary conditions of optimality for Problem 4.2.
Proposition 6.
The minimizer $(\Phi_{t},K_{t})$ of Problem 4.2, or equivalently Problem 4.2, is such that
$\dot{\Phi}_{t}$ $=K_{t}\Phi_{t},$ (21a)
with $\Phi_{0}=I$ and $K_{t}$ is given by (17) and
$R_{t}$ $=-(M_{t}B+P_{t}\Phi_{t}^{\top})N_{t}^{\top}(N_{t}BN_{t}^{\top})^{-1},$ (21b)
where $\text{where }\hskip 10.0pt\dot{P_{t}}$ $=-{K_{t}}^{\top}(K_{t}\Phi_{t}\Sigma_{0}+P_{t}),$ (21c)
and $\text{and }\hskip 10.0ptB$ $=\big{(}\Phi_{t}\Sigma_{0}\Phi_{t}^{\top}+\varepsilon I\big{)}.$ (21d)
Moreover, if $\mathcal{S}\subset\mathrm{GL}(n)$ is the surface defined by
$\mathcal{S}:=\{\Phi\in\mathrm{GL}(n)~{}|~{}\Phi\Sigma_{0}\Phi^{\top}=\Sigma_{1},\,\Phi Y_{0}=Y_{1}\},$ (21e)
then (21a) and (21c) are subject to the mixed boundary conditions $\Phi_{1}\in\mathcal{S}$ and ${P_{1}}\perp T_{\Phi_{1}}\mathcal{S}$, respectively.
Proof 4.4.
As before, we express the control Hamiltonian of Problem 4.2 as
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}K_{t}\Phi_{t}\Sigma_{0}\Phi^{\top}_{t}K_{t}^{\top}+\varepsilon K_{t}K_{t}^{\top}+P_{t}^{\top}\big{(}K_{t}\Phi_{t}\big{)}\big{)},$
for $K_{t}$ satisfying (17) and for $P_{t}$ an $n\times n$ Lagrange multiplier matrix enforcing (3). Equations (21a) and (21c) are obtained respectively from the necessary conditions $\dot{\Phi}_{t}=\partial\mathcal{H}/\partial P_{t}$ and $\dot{P}_{t}=-\partial\mathcal{H}/\partial\Phi_{t}$. Exploiting the form in (17), we rewrite the previous Hamiltonian
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}(M_{t}+R_{t}N_{t})B(M_{t}+R_{t}N_{t})^{\top}\big{)}$
$+{\rm tr}\big{(}P_{t}^{\top}\big{(}M_{t}+R_{t}N_{t}\big{)}\Phi_{t}\big{)},$
for $B$ satisfying (21d). Indeed, direct application of Pontryagin’s maximum principle, $\partial H/\partial R_{t}=0$, results in (21b). Lastly, Equations (10) and (5) define the surface in (21e) and conditions on $\Phi_{1}$ and $P_{1}$ from the proof of Proposition 3 apply verbatim here, which concludes the proof.
5 Examples
Example I
{subcaptionblock}
Figure 1: The gray envelope represents the given $\Sigma_{t\in[0,1]}$ as in (23). The green curve is the resulting trajectory of the tracer between two given endpoints (22).
{subcaptionblock}
Figure 2: Evolution of the entries of the state transition matrix.
Figure 3: Example I
We present an academic example for Problem 1. We assume the positions of a single tracer particle at $t=0,t=1$ are given as
$Y_{0}=\begin{bmatrix}-1\\ 0\end{bmatrix},\ Y_{1}=\begin{bmatrix}0\\ 1\end{bmatrix}.$ (22)
Information on the ensemble is also given as
$\Sigma_{0}=I$ $,\ \Sigma_{1}=\begin{bmatrix}2&\sqrt{2}\\ \sqrt{2}&2\end{bmatrix},$ (23a)
$\Sigma_{t}=\big{(}(1-t)I+$ $t\Sigma_{1}^{\frac{1}{2}}\big{)}\big{(}(1-t)I+t\Sigma_{1}^{\frac{1}{2}}\big{)}^{\top},$ (23b)
for $I$ denoting the identity matrix and $t\in[0,1]$. The flow $\Sigma_{t}$ corresponds to the optimal mass transport between $\Sigma_{0}$ and $\Sigma_{1}$ (McCann displacement interpolation [1]) with no data on tracer particles. This is depicted in Fig. 3, where the gray cylinder represents the evolution of a equi-probability level set of the corresponding Gaussian distributions. The availability of data on the terminal disposition of a tracer particle reveals a rotation component of the flow that cannot be gleaned from the McCann flow in (23b).
Indeed, since $\Phi_{1}{\Phi_{1}}^{\top}=\Sigma_{1}$, $\Phi_{1}=\Sigma_{1}^{\frac{1}{2}}U$, for some $U\in\rm{SO}(2)$, the special orthogonal group of dimension $2$. In the present case where $n=2$, the matrix $U$ is fully determined by the information provided. Therefore, one can solve for $U$ satisfying $\Sigma_{1}^{-\frac{1}{2}}Y_{1}=UY_{0}$, for $Y_{0},Y_{1}$ as in (22). This results in
$U=\begin{bmatrix}\frac{1}{2}\sqrt{2-\sqrt{2}}&\frac{1}{2}\sqrt{\sqrt{2}+2}\\ -\frac{1}{2}\sqrt{\sqrt{2}+2}&\frac{1}{2}\sqrt{2-\sqrt{2}}\end{bmatrix},$
which corresponds to a rotation by $67.5^{\circ}$. Then,
$\Phi_{1}=\begin{bmatrix}0&\sqrt{2}\\ -1&1\end{bmatrix},$ (24)
and the system (13a)-(13b) can be solved as a two-point boundary value problem between $\Phi_{0}=I$ and (24). The numerically obtained state-transitions $\Phi_{t}$ from the shooting method are shown in Fig. 3. These state transitions result in the trajectory depicted in green in Fig. 3, indeed reconciling the information in (22) and (23).
Example II
In contrast to the previous example, we here assume the full trajectory of a single tracer is given as shown in Fig. 6 in green. This is precisely a trajectory between the endpoints
$Y_{0}=\begin{bmatrix}-\frac{\sqrt{3}}{2}\\ \frac{1}{\sqrt{2}}\end{bmatrix},\ Y_{1}=\begin{bmatrix}0\\ -\sqrt{3}\end{bmatrix}.$ (25)
Moreover, we assume the endpoint covariances are
$\Sigma_{0}=I,\ \Sigma_{1}=3I.$
The task here is to numerically demonstrate the solution of Problem 4.2 in the two-dimensional case. Similar to the previous example, the terminal state transition matrix can be fully determined as $\Phi_{1}=\Sigma_{1}^{\frac{1}{2}}O$, for
$O=\begin{bmatrix}-\frac{1}{2}&\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{bmatrix}\in\rm{SO}(2),$
which amounts to a rotation of $-120^{\circ}$. The resulting $\Phi_{1}$ along with $\Phi_{0}$ are used to solve the two-point boundary problem associated with system (21a)-(21d) for $\varepsilon=1$. The obtained entries of $\Phi_{t}$ are presented in Fig. 6. The gray envelope shown in Fig. 6 delineates the equi-probability level sets of the corresponding Gaussian distributions.
{subcaptionblock}
Figure 4: The green curve is the given trajectory of the tracer between two endpoints (25) while the gray envelope depicts the resulting covariances.
{subcaptionblock}
Figure 5: Evolution of the entries of the state transition matrix.
Figure 6: Example II
6 Concluding remarks
The theme of the present work brings up a hitherto unstudied angle in optimal transport problems, where, now, additional data reveal the dynamics of internal degrees of freedom in the flow. Paradigmatically, in either fluid flow or the flow of a flock of birds, tagging specific individuals (tracer particles or birds, respectively) may provide information about rotational components of the collective dynamics that are not reflected in the aggregate distributions.
The work herein explores this theme in the elementary setting of linear dynamics and Gaussian distributions. From a controls perspective, we envision stochastic formulations, where process and measurement noise corrupt the available data, that mirror the deterministic setting presented herein. From an applications standpoint, tracer information is naturally available in several fields. For instance, in computer vision, tracers may correspond to landmarks or facial features, and, in fluid flows, tracers are often injected for precisely the same reasons as postulated herein. In these and other applications, we expect that the conceptualization present in this paper will prove relevant.